OpenStudy (anonymous):

-8x -x + 4x + 7y - 6y + 5

jigglypuff314 (jigglypuff314):

to simplify this know that terms with the same variables can add and subtract with each other like 3x + 2x = 5x so -8x -x + 4x = -9x + 4x = -5x and 7y - 6y = y

jigglypuff314 (jigglypuff314):

like how -8 - 1 + 4 = -9 + 4 = -5 -8x -x + 4x = -9x + 4x = -5x

OpenStudy (anonymous):

Add like terms with like terms X's with X's and Y's with Y's and # with number

OpenStudy (anonymous):

-9x